# Unveiling Latent Causal Rules: A Temporal Point Process Approach for Abnormal Event Explanation

### 03 Feb 2024 v1

@@ 6.2 Healthcare Data Experiments
MIMIC-IV [^1] https://mimic.mit.edu/ is a dataset of electronic health records for patients who have been admitted to the intensive care unit (ICU) [8]. In our study, we focused on 4074 patients diagnosed with sepsis [15], as sepsis is a leading cause of mortality in the ICU, particularly when it progresses to septic shock. Septic shocks are critical medical emergencies, and timely recognition and treatment are crucial for improving survival rates. The objective of this experiment is to identify logic rules associated with septic shocks for both the general population and specific patients. These rules could serve as potential early alarms when abnormal indicators are detected, aiding in timely intervention.
@@
A total of 29 variables associated with sepsis were extracted, including routine vital signs and laboratory values, as suggested and utilized in [9]. We recorded the time points at which these variables first exhibited abnormal values within the 48-hour period preceding the onset of abnormal urine output (i.e., our target event).
Discovered Logic Rules
@@
Real-time urine output was utilized as the health state indicator, as low urine output directly indicates a deteriorate circulatory system and serves as a warning sign for septic shock. Due to the frequent fluctuations in urine output within the ICU setting, only instances where urine output becomes abnormal after maintaining a normal level for at least 48 hours were considered valid target events (i.e., meaningful to predict and explain).
@@
In Tab. 3, we present a portion of the discovered explanatory logic rules and their corresponding learned weights based on our method. These displayed rules have been justified by doctors.
Rule 1 highlights the significance of abnormal arterial blood pressure diastolic as a crucial indicator for septic shock, with a substantial rule weight of 0.566. Therefore, patients should closely pay high attention to their arterial blood pressure diastolic and promptly seek clinical support if it deviates from the normal range.
Rule 2, on the other hand, noted by doctors, may be less relevant as fluctuations in calcium homeostasis can occur in specific patients during sepsis. Rule 3 and Rule 4 highlight the individual significance of abnormal arterial $CO_{2}$ pressure and venous $O_{2}$ pressure as additional early indicators of a compromised circulatory system. These rules bear relatively significant weights of 0.527 and 0.466, respectively.
Rule 5 and Rule 6 demonstrate the combined effects on low urine output. Rule 5, with a weight of 0.464, suggests that when the respiratory rate of patients exceeds 20 breaths per minute and there is a sharp decline in hemoglobin levels, they should be highly vigilant for the possibility of septic shock. Furthermore, Rule 6, which has the highest weight of 0.664, indicates that if the heart rate of patients surpasses 90 beats per minute, blood urea nitrogen (BUN) and white blood cell count (WBC) fall outside the normal ranges, and all three signals occur successively, immediate emergency medical treatment should be sought.
@@
Our methodology incorporated invaluable input from medical experts who thoroughly reviewed our discovered rule results. Their feedback consistently validated the soundness and relevance of our findings. Additionally, we bolstered our discovered rules by referencing and citing supporting evidence from reputable medical sources, with discussions found in Appendix Section 8. We also compared our method with the post-hoc method [3] in terms of interpretability, with the corresponding results in Appendix Section 5.
@@
Figure 6: Rule identification results for patient 14823401. X-axis: time ranges from 0 to 48 h. Y-axis: feature id (29 in total). Low urine output occurred at 48h. Deep blue blocks record the occurrence time of a routine vital sign or abnormal lab value. Learned rule with largest posterior probability (colored in purple): LowUrine $\leftarrow$ Heart Rate $\land$ BUN $\land$ WBC, Heart Rate Before BUN, BUN Before WBC. Other rules with relatively large posterior probability are colored in pink and off white. The darker the color, the greater the posterior probability.
Logic Rule Identification for Each Event
Our method can identify the most likely causal rule specific to individual patients. For instance, let’s consider patient with ID:14823401 as an example and the results are visualized in Fig. 6. Based on the analysis, the rule with the highest posterior probability that the patient is likely to adhere to is Rule 6: LowUrine $\leftarrow$ Heart Rate $\land$ BUN $\land$ WBC. The identified rule will help with the individual healthcare delivery and treatment design.
Event Prediction
We used the same baselines as in the synthetic data experiments, with the MAE serving as the evaluation metric for predicting the occurrence time of “Low Urine” events. Our model’s performance is presented in Tab. 4, where it outperforms all the baselines.
@@
Table 4: Event time prediction MAE on healthcare data.

### 19 Mar 2024 v2

@@ 6.2 Healthcare Data Experiments
MIMIC-IV [^1] https://mimic.mit.edu/ is a dataset of electronic health records for patients who have been admitted to the ICU [15]. In our study, we focused on 4074 patients diagnosed with sepsis [27], as sepsis is a leading cause of mortality in ICU, particularly when it progresses to septic shock. Septic shocks are critical medical emergencies, and timely recognition and treatment are crucial for improving survival rates. The objective of this experiment is to identify logic rules associated with septic shocks for both the general population and specific patients. These rules could serve as potential early alarms when abnormal indicators are detected, aiding in timely intervention.
@@
Routine Vital Signs and Lab Values A total of 29 variables associated with sepsis were extracted, including routine vital signs and laboratory values, as suggested and utilized by [16]. Detailed illustration can be found in Appendix Sec.F. We recorded the time points at which these variables first exhibited abnormal values within the 48 hours preceding the onset of abnormal urine output (i.e., our target event).
@@
Discovered Logic Rules Real-time urine output was utilized as the health state indicator, as low urine output directly indicates a deteriorated circulatory system and serves as a warning sign for septic shock. Due to the frequent fluctuations in urine output within the ICU setting, only instances where urine output becomes abnormal after maintaining a normal level for at least 48 hours were considered valid target events (i.e., meaningful to predict and explain).
@@
In Tab.3, we present a portion of the discovered explanatory logic rules and their corresponding learned weights. These displayed rules have been justified by doctors or supported by relevant literature, demonstrating their clinical significance.
@@
Our methodology incorporated invaluable viewpoints from experts who thoroughly reviewed our discovered rules. Their feedback consistently validated the soundness and relevance of our findings. Additionally, we bolstered our discovered rules by referencing supporting evidence from reputable medical sources, with detailed discussions and explanations found in Appendix Sec.G.
We also compared our model with the post-hoc method [6] in terms of interpretability, with the corresponding results in Appendix Sec.D. While the post-hoc methods are easier to implement, our model surpasses them in terms of accuracy and interpretability. Moreover, our model remains reliable even in small data scenarios, which is challenging for post-hoc methods relying on black-box models to achieve.
@@
Figure 6: Rule identification results for patient 14823401. X-axis: time ranges from 0 to 48h. Y-axis: feature ID (29 in total). Low urine output occurred at 48h. Deep blue blocks record the occurrence time of a routine vital sign or abnormal lab value. Identified rule with largest posterior probability (colored in purple), Rule 6: LowUrine $\leftarrow$ Heart Rate $\land$ BUN $\land$ WBC, Heart Rate Before BUN, BUN Before WBC. Other rules with relatively large posterior probability are colored in pink and off white. The darker the color, the greater the posterior probability.
Logic Rule Identification for Each Event Our method can identify the most likely causal rule specific to individual patients. For instance, let’s consider the patient with ID 14823401 as an example, and the results are visualized in Fig.6. Based on the analysis, the rule with the highest posterior probability that the patient is likely to adhere to is Rule 6: LowUrine $\leftarrow$ Heart Rate $\land$ BUN $\land$ WBC, indicating that this patient requires closer monitoring of Heart Rate, BUN, and WBC levels exceeding the normal range in their daily life, which likely serves as an early warning sign. This example sufficiently illustrates the effectiveness of our method at the individual patient level, showcasing its potential for individual healthcare delivery and treatment design. Event Prediction We used the same baselines as in the synthetic data experiments, with the MAE serving as the evaluation metric for predicting the occurrence time of “Low Urine” events. Our model’s performance is presented in Tab.4, from which one can see that our model outperforms all other baselines.
@@
Table 4: Event time prediction MAE (unit/hour) on healthcare data.